Calculate the molality of a 35.0% (by mass) aqueous solution of sulfuric acid (H2SO4).

(Molar mass of H2SO4 is 98.02 g/mol and for H2O is 18.02 g/mol)
Answer #1

Suppose we have 100 grams of solution

Mass of H2SO4 = 35 grams

Mass of water = 100 - 35 = 65 grams

Moles of H2SO4 = mass / molar mass

= 35 / 98.02

= 0.357

Mass of solvent = 65 grams or 0.065 Kg

Molality = moles / mass of solvent ( in kg)

= 0.357 / 0.065

= 5.492

If you have any questions please comment

If you satisfied with the solution please rate it thanks